Call us: +44 (0)330 043 1625
All systems are operational

Past Incidents

Monday 29th July 2013

No incidents reported

Sunday 28th July 2013

No incidents reported

Saturday 27th July 2013

No incidents reported

Friday 26th July 2013

No incidents reported

Thursday 25th July 2013

No incidents reported

Wednesday 24th July 2013

Network Packet loss

We have become aware of a spike of traffic from external sources.

We have null routed the traffic at our network edge and expect the problem to be resolved shortly.

  • Update (11:45): The high levels of inbound traffic is continuing over different IP blocks
  • Update (12:00): The issue is still being investigated
  • Update (12:15): The issue is still being investigated
  • Update (12:45): The affected IP blocks have been de-announced at our main peering points and via all carriers, at present the BGP route announcement is propagating
  • Update (13:00): The affected carriers are now no longer carrying the dDOS traffic and all services are restored to 100%

Tuesday 23rd July 2013

No incidents reported